OpenStudy (anonymous):

What is the solution set of this system of equations? y = x^2 − 3x − 4 x = y + 8 A) {(-1,0), (4,0)} B) {(8,0), (0,8)} C) {(0,0)} D) {(2,-6)} E) No real solution

OpenStudy (anonymous):

The answer set is x = 2, y = -6. Refer to WolframAlpha here: http://www.wolframalpha.com/input/?i=y+%3D+x%5E2+-+3x+-4+and+y+%3D+x-8

OpenStudy (anonymous):

plug in x=y+8 in the first one

OpenStudy (jdoe0001):

\(\bf y = x^2 - 3x - 4\\ x = y + 8 \\ \quad \\ 0 = x^2 - 3x - 4\quad \textit{factor the equation to find "x"}\)